[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed May 11, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Wed May 11 14:09:41 UTC 2016


Here is today's summary of NHC computer operations:
--- Monica

----------

TAFB:
None
HSU:
None
Other items:
- WCOSS production switch from tide to gyre occurred at 12Z this morning.  
Therefore, forecasters will have access to the new GFS model starting  
today. TSB has confirmed that NO CHANGES are needed in forecaster SPFs, and  
GFE will not be affected either.
- SSU, with support from TSB and HSU, continues to work with NCO this week  
to test Psurge and the HSOFS model. Tests will be done at 15Z and 21Z every  
day this week.
NHC RFCs:
None

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160511/8fd451a0/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list